add and subtract the following
1) (21x^2y+4xy^2-7x)-(8x^2y+11xy^2-3x)
2) (8x^2+11xy-17y^2)+(2x^2+4xy-18y^2)
3) (-4x^2y-2z^2y)-(8x^2y-4z^2y)+(-6x^2y+14xy^2)
4) (2/3x^3-1/2x+1/4)-(2/5x^2-1/3x+1/9)
qustion 1 I got 13x^2y-15xy^2-10x
question 2 I got 10x^2+15xy+35y^2
question 3 I got -6x^2y+-6z^2y+14xy^2
question 4 I got 2/3x^3-2/5x^2-1/6x+1/13
are these correct

Answer by ankor@dixie-net.com(22740)

Add and subtract the following
1) (21x^2y+4xy^2-7x)-(8x^2y+11xy^2-3x)
you have a neg outside the bracket you must change signs inside the br
It would be like this:
21x^2y + 4xy^2 - 7x - 8x^2y - 11xy^2 + 3x; combine these and you have:
13x^2 - 7xy^2 - 4x
:
2) (8x^2+11xy-17y^2)+(2x^2+4xy-18y^2)
In this one everything is right except the last sign should be -35y^2
:
3) (-4x^2y-2z^2y)-(8x^2y-4z^2y)+(-6x^2y+14xy^2)
After removing the brackets you have:
-4x^2y - 2z^2y - 8x^2y + 4z^2y -6x^2y + 14xy^2 =
-18x^2y + 2z^2y + 14xy^2
:
4) (2/3x^3-1/2x+1/4)-(2/5x^2-1/3x+1/9)
After removing the brackets
(2/3)x^3 - (1/2)x + 1/4 -(2/5)x^2 + (1/3)x - 1/9 =
(2/3)x^3 - (2/5)x^2 - (1/6)x + 5/36; you were right except for the last term